OpenStudy (anonymous):

cos(3x)+0.5=0 how many solutions are there.

OpenStudy (anonymous):

@satellite73 can u help me?

OpenStudy (anonymous):

@zarkon can u help me?

OpenStudy (anonymous):

there are an infinite number of solutions because cosine is periodic. start with \[\cos(3x)=-\frac{1}{2}\] making \[3x=\frac{\pi}{3}+2k\pi\] or \[3x=\frac{2\pi}{3}2k\pi\] and then divide by 3 to solve for x